OpenStudy (anonymous):

Leila is a software saleswoman. Her base salary is , and she makes an additional for every copy of English is Fun she sells. Let represent her total pay (in dollars), and let represent the number of copies of English is Fun she sells. Write an equation relating to . Then use this equation to find her total pay if she sells copies of English is Fun.

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

OK, you can just make up any number for her base salary you want. So I'm going to say her base salary is $30,000. You can set her commission for selling each copy of the software at any number you want. So I'm going to say she makes $1.50 for every copy of software she sells. I'm going use the variable p for pay and n for the number copies sold. My sales to total pay equation will look like this: 30,000 + 1.50n = p If she sells 1,000 copies of the software, her total pay is: 30,000 + 1.50

OpenStudy (anonymous):

30,000+1.50(1000)= p 30,000 +1,500 = 31,500 Her total pay is $31,500